# Optimal Aggregation of Prediction Intervals under Unsupervised Domain Shift

### v1

@@ Abstract
As machine learning models are increasingly deployed in dynamic environments, it becomes paramount to assess and quantify uncertainties associated with distribution shifts. A distribution shift occurs when the underlying data-generating process changes, leading to a deviation in the model’s performance. The prediction interval, which captures the range of likely outcomes for a given prediction, serves as a crucial tool for characterizing uncertainties induced by their underlying distribution. In this paper, we propose methodologies for aggregating prediction intervals to obtain one with minimal width and adequate coverage on the target domain under unsupervised domain shift, under which we have labeled samples from a related source domain and unlabeled covariates from the target domain. Our analysis encompasses scenarios where the source and the target domain are related via i) a bounded density ratio, and ii) a measure-preserving transformation. Our proposed methodologies are computationally efficient and easy to implement. Beyond illustrating the performance of our method through a real-world dataset, we also delve into the theoretical details. This includes establishing rigorous theoretical guarantees, coupled with finite sample bounds, regarding the coverage and width of our prediction intervals. Our approach excels in practical applications and is underpinned by a solid theoretical framework, ensuring its reliability and effectiveness across diverse contexts [^1] For the code to reproduce the results in our paper, see here..
@@ 1 Introduction
In the modern era of big data and complex machine learning models, extensive data collected from diverse sources are often used to build a predictive model. However, the assumption of independent and identically distributed (i.i.d.) data is frequently violated in practical scenarios. Take algorithmic fairness as an example: historical data often exhibit sampling biases towards certain groups, like females being underrepresented in credit card data. Over time, the differences in group proportions have diminished, leading to distribution shifts. Consequently, models trained on historical data may face shifted distributions during testing, and proper adjustments is needed. Distribution shift has garnered significant attention from statistical and machine learning communities under various names, i.e., transfer learning (Pan and Yang,, 2009; Weiss et al.,, 2016), domain adaptation (Farahani et al.,, 2021), domain generalization (Zhou et al.,, 2022; Wang et al.,, 2022), continual learning (De Lange et al.,, 2021; Mai et al.,, 2022), multitask learning (Zhang and Yang,, 2021) etc. While numerous methods are available in the literature for training predictive models under distribution shift, uncertainty quantification under distribution shift has received relatively scant attention despite its crucial importance. One notable exception is conformal prediction under distribution shift; Tibshirani et al., (2019) proposed a variant of standard conformal inference methods to accommodate test data from a distinct distribution from the training data under the covariate shift. Recently, Gibbs and Candes, (2021) introduced an adaptive conformal inference approach suitable for continuously changing distributions over time. Additionally, quantile regression under distribution shift offers another avenue for addressing uncertainty quantification under distribution shift (Eastwood et al.,, 2022).
@@
Although few methods exist for constructing prediction intervals under distribution shift, most focus primarily on ensuring coverage guarantee rather than minimizing interval width. This prompts the immediate question: Can we generate prediction intervals in the target domain that provide both i) coverage guarantee and ii) minimal width? This paper seeks to address this question by leveraging model aggregation techniques. Suppose we have $K$ different methods for constructing prediction intervals in the source domain. Our proposed approach efficiently combines these methods to produce prediction intervals in the target domain with adequate coverage and minimal width. When individual methods are the elementary basis functions, such as the kernel basis, the resulting aggregation is indeed a construction of the prediction interval based on the basis functions. Our methodology draws inspiration primarily from recent work (Fan et al.,, 2023) on prediction interval aggregation under the i.i.d. setting. However, a key distinction lies in our focus on unsupervised domain adaptation, where we can access labeled samples from the source and unlabeled samples from the target domain. Certain assumptions regarding the similarities between these domains are necessary to facilitate knowledge transfer from the source to the target domain. We explore two types of similarities in this paper: i) covariate shift, where we assume that the distribution of the response variable $Y$ given $X$ is consistent across both domains, albeit the distribution of $X$ may differ, and ii) domain shift, where we assume that the conditional distribution of $Y$ given $X$ remains unchanged up to a measure-preserving transformation. Covariate shift is a well-explored concept in transfer learning and has also garnered attention in uncertainty quantification. It allows different distributions of $X$ while maintaining identical conditional distributions $Y|X$ across domains. For constructing conformal prediction intervals within this framework, see Tibshirani et al., (2019); Hu and Lei, (2023); Yang et al., (2022); Lei and Candès, (2021) and references therein. On the other hand, distribution shift is more general, allowing both the distribution of $X$ and the conditional distribution of $Y|X$ to differ across domains. Our methods in this context draw upon domain matching principles via transport map, as proposed in Courty et al., (2014) and further elaborated in subsequent works like Courty et al., (2016, 2017); Redko et al., (2017), among others. The key assumption is the existence of a measure-preserving/domain-aligning map $T$ from the target to the source domain, such that the conditional distribution of $Y|X$ on the target domain matches $Y|T(X)$ on the source domain, i.e., conditional distributions matches upon domain alignment. The case where the domain-aligning map is the optimal transport map has received considerable attention in the literature, e.g., see Courty et al., (2014, 2016, 2017); Xu et al., (2020). Empirical evidence supports the efficacy of domain alignment through optimal transport maps across various datasets. For instance, in Xu et al., (2020), a variant of this method is applied for domain adaptation in image recognition tasks, such as recognizing similarities between USPS (Hull,, 1994), MNIST (LeCun et al.,, 1998), and SVHN digit images (Netzer et al.,, 2011), as well as between different types of images in the Office-home dataset (Venkateswara et al.,, 2017), including artistic and product images. Additionally, in Courty et al., (2014), the authors explore the impact of domain alignment via optimal transport maps on the face recognition problem, where different poses give rise to distinct domains. However, most of these works concentrate on training predictors that perform well on the target domain without any guarantee regarding uncertainty quantification. To our knowledge, this is the first work to propose a method with rigorous theoretical guarantees for constructing prediction intervals on the target domain under the domain-aligning assumption within an unsupervised domain adaptation framework. We now summarize our contributions.
@@
Our Contributions: This paper introduces a novel methodology for aggregating various prediction methods available on the source domain to construct a unified prediction interval on the target domain under both covariate shift and domain shift assumptions. Our approach is simple and easy to implement and requires solving a convex optimization problem, which can even be simplified to a linear program problem in certain scenarios. We also establish rigorous theoretical guarantees, presenting finite sample concentration bounds to demonstrate that our method achieves adequate coverage with a small width. Furthermore, our methodology extends beyond model aggregation; it can be used to construct efficient prediction intervals from any convex collection of candidate functions. In the paper, we adopt this broader perspective, discussing how the aggregation of prediction intervals emerges as a particular case. Lastly, we validate the effectiveness of our approach by analyzing a real-world dataset.
@@ 2.1 Problem formulation
$\min_{f\in\mathcal{F}}\quad\ \mathbb{E}_{T}[u(X)-l(X)],\ \ \textrm{s.t.}\ \ \mathbb{P}_{T}\left(l(X)\leq Y\leq u(X)\right)\geq 1-\alpha\,.$ (2.1)
@@ 3 Covariate shift with bounded density ratio
In this section, we focus on the covariate shift problems, where the marginal densities $p_{S}(X)$ and $p_{T}(X)$ of the covariates may vary between the source and target domains, albeit the conditional distribution $Y|X$ remains same. Denote by $m_{0}(x)=\mathbb{E}_{T}[Y|X=x]=\mathbb{E}_{S}[Y|X=x]$, the conditional mean function. For the ease of the presentation, we assume $m_{0}$ is known. If unknown, one may use the labeled source data to estimate it using a suitable parametric/non-parametric estimate (e.g., splines, local polynomial, or deep neural networks), subsequently substituting $m_{0}$ with $\hat{m}$ in our approach. The density ratio of the source and the target distribution of $X$ is denoted by $w_{0}(x):=p_{T}(x)/p_{S}(x)$. We henceforth assume that the density ratio is uniformly bounded:
@@
Step 1: (Shape estimation) Obtain an initial estimate $\hat{f}_{\mathrm{init}}$ via by solving (3.1) for $\alpha=0$ (to capture the shape):
@@ 5 Application
In this section, we illustrate the effectiveness of our method using the airfoil dataset from the UCI Machine Learning Repository (Dua and Graff,, 2019). This dataset includes $1503$ observations, featuring a response variable $Y$ (scaled sound pressure level) and a five-dimensional covariate $X$ (log of frequency, angle of attack, chord length, free-stream velocity, log of suction side displacement thickness). We assess and compare the performance of our prediction intervals in terms of coverage and width with those generated by the weighted split conformal prediction method described in Tibshirani et al., (2019).
We use the same data-generating process described in Tibshirani et al., (2019) to facilitate a direct comparison. We have run experiments 200 times; each time, we randomly partitioned the data into two parts $\mathcal{D}_{\rm train}$ and $\mathcal{D}_{\rm test}$, where $\mathcal{D}_{\rm train}$ contains $75\%$ of the data, and $\mathcal{D}_{\rm test}$ contains $25\%$ of the data. Following Tibshirani et al., (2019), we shift the distribution of the covariates of $\mathcal{D}_{\rm test}$ by weighted sampling with replacement, where the weights are proportional to
@@
We now present the histograms of the coverage and the average bandwidth of our method, and a more general version of weighted conformal prediction in Tibshirani et al., (2019) over $200$ experiments (see Appendix B for details),
@@ Appendix A Proofs
Finally as $f-f^{*}$ is upper bounded by $F^{\prime}=B_{\mathcal{F}}+\|f^{*}\|_{\infty}$ (as $f$ is uniformly upper bounded by F). Therefore, by Mcdiarmid’s inequality, we with have with probability $1-e^{t}$:
@@
We next apply vector-valued Leduox-Talagrand contraction inequality on the function $\phi$ (equation (1) of Maurer, (2016)), to obtain the following bound on the Rademacher complexity:
@@
where the first inequality used $\ell_{\delta}(x)\geq\mathds{1}(x\leq-\delta)$, second inequality uses the fact that sample average of $\ell_{\delta}$ over $\mathcal{D}_{S,1}$ is 00 by the definition of $\hat{f}_{\mathrm{init}}$, third inequality uses Leduox-Talagrand contraction inequality observing that $\ell_{\delta}$ is $1/\delta$-Lipschitz. This completes the proof.

### v2

@@ Abstract
As machine learning models are increasingly deployed in dynamic environments, it becomes paramount to assess and quantify uncertainties associated with distribution shifts. A distribution shift occurs when the underlying data-generating process changes, leading to a deviation in the model’s performance. The prediction interval, which captures the range of likely outcomes for a given prediction, serves as a crucial tool for characterizing uncertainties induced by their underlying distribution. In this paper, we propose methodologies for aggregating prediction intervals to obtain one with minimal width and adequate coverage on the target domain under unsupervised domain shift, under which we have labeled samples from a related source domain and unlabeled covariates from the target domain. Our analysis encompasses scenarios where the source and the target domain are related via i) a bounded density ratio, and ii) a measure-preserving transformation. Our proposed methodologies are computationally efficient and easy to implement. Beyond illustrating the performance of our method through real-world datasets, we also delve into the theoretical details. This includes establishing rigorous theoretical guarantees, coupled with finite sample bounds, regarding the coverage and width of our prediction intervals. Our approach excels in practical applications and is underpinned by a solid theoretical framework, ensuring its reliability and effectiveness across diverse contexts.
@@ 1 Introduction
In the modern era of big data and complex machine learning models, extensive data collected from diverse sources are often used to build a predictive model. However, the assumption of independent and identically distributed (i.i.d.) data is frequently violated in practical scenarios. Take algorithmic fairness as an example: historical data often exhibit sampling biases towards certain groups, like females being underrepresented in credit card data. Over time, the differences in group proportions have diminished, leading to distribution shifts. Consequently, models trained on historical data may face shifted distributions during testing, and proper adjustments are needed. Distribution shift has garnered significant attention from statistical and machine learning communities under various names, i.e., transfer learning (Pan and Yang,, 2009; Weiss et al.,, 2016), domain adaptation (Farahani et al.,, 2021), domain generalization (Zhou et al.,, 2022; Wang et al.,, 2022), continual learning (De Lange et al.,, 2021; Mai et al.,, 2022), multitask learning (Zhang and Yang,, 2021) etc. While numerous methods are available in the literature for training predictive models under distribution shift, uncertainty quantification under distribution shift has received relatively scant attention despite its crucial importance. One notable exception is conformal prediction under distribution shift; Tibshirani et al., (2019) proposed a variant of standard conformal inference methods to accommodate test data from a distinct distribution from the training data under the covariate shift. Recently, Gibbs and Candes, (2021) introduced an adaptive conformal inference approach suitable for continuously changing distributions over time. Additionally, quantile regression under distribution shift offers another avenue for addressing uncertainty quantification under distribution shift (Eastwood et al.,, 2022).
Although few methods exist for constructing prediction intervals under distribution shift, most focus primarily on ensuring coverage guarantee rather than minimizing interval width. This prompts the immediate question:
Can we generate prediction intervals in the target domain that provide both i) coverage guarantee and ii) minimal width?
@@
This paper seeks to address this question by leveraging model aggregation techniques (Nowotarski and Weron,, 2015; Maciejowska et al.,, 2016; Carlsson et al.,, 2014; Vovk,, 2015; Hosen et al.,, 2014). Suppose we have $K$ different methods for constructing prediction intervals in the source domain. Our proposed approach efficiently combines these methods to produce prediction intervals in the target domain with adequate coverage and minimal width. When individual methods are the elementary basis functions, such as the kernel basis, the resulting aggregation is indeed a construction of the prediction interval based on the basis functions. Our methodology draws inspiration primarily from recent work (Fan et al.,, 2023) on prediction interval aggregation under the i.i.d. setting. However, a key distinction lies in our focus on unsupervised domain adaptation, where we can access labeled samples from the source and unlabeled samples from the target domain. Certain assumptions regarding the similarities between these domains are necessary to facilitate knowledge transfer from the source to the target domain. We explore two types of similarities in this paper: i) covariate shift, where we assume that the distribution of the response variable $Y$ given $X$ is consistent across both domains, albeit the distribution of $X$ may differ, and ii) domain shift, where we assume that the conditional distribution of $Y$ given $X$ remains unchanged up to a measure-preserving transformation. Covariate shift is a well-explored concept in transfer learning and has also garnered attention in uncertainty quantification. It allows different distributions of $X$ while maintaining identical conditional distributions $Y|X$ across domains. For constructing conformal prediction intervals within this framework, see Tibshirani et al., (2019); Hu and Lei, (2023); Yang et al., (2022); Lei and Candès, (2021) and references therein. On the other hand, distribution shift is more general, allowing both the distribution of $X$ and the conditional distribution of $Y|X$ to differ across domains. Our methods in this context draw upon domain matching principles via transport map, as proposed in Courty et al., (2014) and further elaborated in subsequent works like Courty et al., (2016, 2017); Redko et al., (2017), among others. The key assumption is the existence of a measure-preserving/domain-aligning map $T$ from the target to the source domain, such that the conditional distribution of $Y|X$ on the target domain matches $Y|T(X)$ on the source domain, i.e., conditional distributions matches upon domain alignment. The case where the domain-aligning map is the optimal transport map has received considerable attention in the literature, e.g., see Courty et al., (2014, 2016, 2017); Xu et al., (2020). Empirical evidence supports the efficacy of domain alignment through optimal transport maps across various datasets. For instance, in Xu et al., (2020), a variant of this method is applied for domain adaptation in image recognition tasks, such as recognizing similarities between USPS (Hull,, 1994), MNIST (LeCun et al.,, 1998), and SVHN digit images (Netzer et al.,, 2011), as well as between different types of images in the Office-home dataset (Venkateswara et al.,, 2017), including artistic and product images. Additionally, in Courty et al., (2014), the authors explore the impact of domain alignment via optimal transport maps on the face recognition problem, where different poses give rise to distinct domains. However, most of these works concentrate on training predictors that perform well on the target domain without any guarantee regarding uncertainty quantification. To our knowledge, this is the first work to propose a method with rigorous theoretical guarantees for constructing prediction intervals on the target domain under the domain-aligning assumption within an unsupervised domain adaptation framework. We now summarize our contributions.
@@
Our Contributions: This paper introduces a novel methodology for aggregating various prediction methods available on the source domain to construct a unified prediction interval on the target domain under both covariate shift and domain shift assumptions. Our approach is simple and easy to implement and requires solving a convex optimization problem, which can even be simplified to a linear program problem in certain scenarios. We also establish rigorous theoretical guarantees, presenting finite sample concentration bounds to demonstrate that our method achieves adequate coverage with a small width. Furthermore, our methodology extends beyond model aggregation; it can be used to construct efficient prediction intervals from any convex collection of candidate functions. In the paper, we adopt this broader perspective, discussing how the aggregation of prediction intervals emerges as a particular case. Lastly, we validate the effectiveness of our approach by analyzing real-world datasets.
We also want to highlight the differences between our method and a related method proposed in (Fan et al.,, 2023). We deal with unsupervised domain adaptation, i.e., we do not observe any label from the target domain, in contrast to (Fan et al.,, 2023), which only deals with i.i.d. data. Hence, significant changes in methodology are required to address the domain shift. Furthermore, as pointed out in Section 3, the shift may cause the optimization problem non-convex, for which we need to introduce a convex surrogate (e.g., the hinge function), leading to additional theoretical challenges.
@@ 2.1 Problem formulation
$\min_{u,l}\quad\ \mathbb{E}_{T}[u(X)-l(X)],\ \ \textrm{s.t.}\ \ \mathbb{P}_{T}\left(l(X)\leq Y\leq u(X)\right)\geq 1-\alpha\,.$ (2.1)
@@ 3 Covariate shift with bounded density ratio
In this section, we focus on the covariate shift problems, where the marginal densities $p_{S}(X)$ and $p_{T}(X)$ of the covariates may vary between the source and target domains, albeit the conditional distribution $Y|X$ remains the same. Denote by $m_{0}(x)=\mathbb{E}_{T}[Y|X=x]=\mathbb{E}_{S}[Y|X=x]$, the conditional mean function. For the ease of the presentation, we assume $m_{0}$ is known. If unknown, one may use the labeled source data to estimate it using a suitable parametric/non-parametric estimate (e.g., splines, local polynomial, or deep neural networks), subsequently substituting $m_{0}$ with $\hat{m}$ in our approach. The density ratio of the source and the target distribution of $X$ is denoted by $w_{0}(x):=p_{T}(x)/p_{S}(x)$. We henceforth assume that the density ratio is uniformly bounded:
@@
Step 1: (Shape estimation) Obtain an initial estimate $\hat{f}_{\mathrm{init}}$ by solving (3.1) for $\alpha=0$ (to capture the shape):
@@ 5 Application
In this section, we illustrate the effectiveness of our method by applying it to five different datasets: i) airfoil dataset (Dua and Graff,, 2019), ii) real estate data (Yeh,, 2018), iii) energy efficiency data (Tsanas and Xifara, 2012b, ), iv) appliance energy prediction data (Candanedo,, 2017), and v) ET Dataset (ETT-small) (Zhou et al.,, 2021). The first four datasets are freely available in the UCI repository, and the last dataset can be found in this GitHub link. Here, we illustrate the procedure using the airfoil dataset, and the details of our experiments using the other datasets can be found in Appendix C. The airfoil dataset includes $1503$ observations, featuring a response variable $Y$ (scaled sound pressure level) and a five-dimensional covariate $X$ (log of frequency, angle of attack, chord length, free-stream velocity, log of suction side displacement thickness). We assess and compare the performance of our prediction intervals in terms of coverage and width with those generated by the weighted split conformal prediction method described in Tibshirani et al., (2019). We use the same data-generating process described in Tibshirani et al., (2019) to facilitate a direct comparison. We run experiments 200 times; each time, we randomly partition the data into two parts $\mathcal{D}_{\rm train}$ and $\mathcal{D}_{\rm test}$, where $\mathcal{D}_{\rm train}$ contains $75\%$ of the data, and $\mathcal{D}_{\rm test}$ contains $25\%$ of the data. Following Tibshirani et al., (2019), we shift the distribution of the covariates of $\mathcal{D}_{\rm test}$ by weighted sampling with replacement, where the weights are proportional to
@@
In Figure 1, we present the histograms of the coverage and the average bandwidth of our method, and a more general version of weighted conformal prediction in Tibshirani et al., (2019) over $200$ experiments (see Appendix B for details),
@@
Figure 1: Experiments on Airfoil data using Algorithm 1
@@
5.1 Robustness of our method
One of the strengths of our approach lies in its resilience to the misspecification of certain components. The core idea behind our method is to combine multiple predictors to create a prediction interval that ensures sufficient coverage while maintaining a narrow average width. These individual prediction intervals may be based on estimators of conditional quantiles, means, variances, and other metrics. If some of the component prediction intervals exhibit poor performance, whether due to inadequate coverage or excessive width, our method typically assigns them lower weights, making it robust to their deficiencies. In contrast, other conformal methods that heavily depend on a single component tend to underperform, particularly with respect to average width. To illustrate this phenomenon, we evaluate our method under model misspecification using a simple one-dimensional simulation setup.
$X\sim{\sf unif}([-1,1]),\ \ \xi\sim{\sf unif}([-1,1]),\ \ X\perp\!\!\!\perp\xi$
$Y=\sqrt{1+25X^{4}}\xi\,.$
As mentioned in the previous subsection, our method aggregates six predictor intervals, including an estimator for the conditional variance function (Estimator 5). The weighted variance-adjusted conformal prediction interval (WVAC) relies on accurately estimating this conditional variance. We estimate the conditional variance using a random forest with varying depths ({3, 5, 7, 15}). For simulation purpose, we generate $n=2500$ samples, keeping $75\%$ as source data and resampling the remaining 25% with weighted samples proportional to $w(x)\propto(1+{\sf exp}{(-2x)})^{-1}$. As depth increases, overfitting leads to poor out-of-sample variance predictions. Table 1 summarizes our findings over 100 Monte Carlo iterations, showing that WVAC’s average width increases with depth, while our method’s average width remains stable.
Max depth Avg. width–Our Method Avg. width–WVAC
3 2.07(0.975) 3.08(0.9712)
5 2.07(0.95) 3.28(0.9664)
7 2.068(0.94) 3.33(0.97)
15 2.08(0.97) 5.00(0.97)
Table 1: Robustness of our method and WVAC. The number inside the parenthesis is the median of coverage over these Monte Carlo iterations.
@@ Appendix A Proofs
Finally as $f-f^{*}$ is upper bounded by $F^{\prime}=B_{\mathcal{F}}+\|f^{*}\|_{\infty}$ (as $f$ is uniformly upper bounded by F). Therefore, by Mcdiarmid’s inequality, we have with probability $1-e^{-t}$:
@@
We next apply vector-valued Ledoux-Talagrand contraction inequality on the function $\phi$ (equation (1) of Maurer, (2016)), to obtain the following bound on the Rademacher complexity:
@@
where the first inequality used $\ell_{\delta}(x)\geq\mathds{1}(x\leq-\delta)$, second inequality uses the fact that sample average of $\ell_{\delta}$ over $\mathcal{D}_{S,1}$ is 00 by the definition of $\hat{f}_{\mathrm{init}}$, third inequality uses Ledoux-Talagrand contraction inequality observing that $\ell_{\delta}$ is $1/\delta$-Lipschitz. This completes the proof.
@@ B.3 Boxplots to compare coverage and bandwidth
Appendix C Additional Experiments
In this section, we present the details of the additional experiments based on four other datasets, as mentioned in Section 5. We compare our methodology against two other widely used conformal methods, namely Weighted Variance-adjusted Conformal (WVAC) (Tibshirani et al.,, 2019), and Weighted Quantile-adjusted Conformal (WQC)—which is a variant of WVAC, where the score function is changed to the conditional quantile Romano et al., (2019).
C.1 Experiment 1: Real estate data
The Real Estate Valuation dataset available at the UCI Machine Learning Repository contains data used to estimate the value of real estate properties. Collected from Taipei, Taiwan, this dataset is useful for predictive modeling in the real estate sector. It consists of 414 instances, with each entry representing a different real estate transaction. This dataset was originally introduced in Tsanas and Xifara, 2012a . In this dataset, the goal is to predict the house price per unit area based on the 6 other features, namely i) transaction date, ii) house age, iii) distance to the nearest MRT station, iv) number of convenience stores, v) latitude and vi) longitude. The construction of shifted data (with $\beta=(-1,0,-1,0,1,1)$) and implementation procedure are the same as in Section 5. Table 2 and Figure 4 presented the results over 200 Monte Carlo iterations. It is evident from the table that our method produced a small average width in comparison to the other methods while maintaining the coverage guarantee.
Real Estate Data
Outcome Our Method WVAC WQC
Coverage (Median) 0.98 0.962 0.971
Coverage (IQR) 0.058 0.048 0.048
Bandwidth (Median) 36.889 46.392 46.858
Bandwidth (IQR) 15.001 19.027 14.189
Bandwidth (Median for Coverage $>$ 95%) 40.774 50.703 51.031
Table 2: Experimental results for the real estate data
C.2 Experiment 2: Energy efficiency data
The Energy Efficiency dataset available at the UCI Machine Learning Repository is designed to help predict the heating load and cooling load requirements of buildings. The dataset was originally introduced in Yeh and Hsu, (2018). The dataset includes various building parameters such as: i) wall area, ii) surface area, iii) roof area, iv) orientation, etc. The goal is to predict the heating load based on 8 other covariates. The construction of shifted data (with $\beta=(-1,0,1,0,-1,0,0,-1)$) and implementation procedure are the same as in Section 5. Our results over 200 Monte Carlo iterations are presented in Table 3 and Figure 4. The table shows that our method produced a smaller bandwidth than WVAC. While WQC has a smaller median bandwidth, it sacrifices coverage. The last row indicates that for experiments with coverage $\geq 95\%$, WQC’s median average width is significantly larger than ours. Thus, whenever WQC provides adequate coverage, its bandwidth is much larger than ours.
Energy efficiency data
Outcome Our Method WVAC WQC
Coverage (Median) 0.995 0.969 0.973
Coverage (IQR) 0.047 0.036 0.05
Bandwidth (Median) 4.332 5.045 2.842
Bandwidth (IQR) 1.358 3.269 2.551
Bandwidth (Median for Coverage $>$ 95%) 4.373 5.681 4.94
Table 3: Experimental results for the energy efficiency data
C.3 Experiment 3: Appliances Energy Prediction Dataset
Appliances Energy Prediction Dataset is freely available from the UCI repository. This dataset is a time series data with 28 covariates and one response variable. We used data from 2016-01-11 to 2016-02-15 (5000 samples) as our training set and data from 2016-05-13 to 2016-05-27 (2000 samples) as our testing set. Since the source and target data are from different time periods, this experiment involves a non-synthetic real-world time shift. The results based on our Algorithm 2 are presented below:
Appliances Energy Prediction Dataset
Outcome Our Method WVAC WQC
Coverage 0.95 1.00 1.00
Bandwidth 461.69 6809.87 2032.12
Table 4: Experimental results for the Appliances Energy Prediction Dataset
C.4 Experiment 4: ETDataset
We applied our method to the ETDataset (ETT-small) from (Zhou et al.,, 2021), which contains hourly-level data from two electricity transformers at two different stations, including load and oil temperature measurements. Each data point consists of 8 features, including the date of the point, the predictive value ”oil temperature”, and 6 different types of external power load features. For our experiment, we used the data from one transformer during the period from July 1, 2016, to November 2, 2016, as our source data, and data from the same time period from the other transformer as our target data. As our source data and the target data are from different locations, we have a geographical covariate shift; see (Zhou et al.,, 2021) for more details. Our results are as follows:
ETDataset
Outcome Our Method WVAC WQC
Coverage 0.976 0.982 0.842
Bandwidth 41.525 57.9 54.981
Table 5: Experimental results for the ETDataset
C.5 Experiment 5: Airfoil data
In Section 5, we have implemented our Algorithm 1 on the Airfoil data (Tibshirani et al.,, 2019) to showcase the efficacy of our method. Here, additionally, we implement Algorithm 2 on the same dataset to evaluate the performance of our second method based on optimal transport-based domain alignment. Here, the data shifting procedure (to create data from the target domain with a shifted distribution) is different from the rest of the experiments; we use the linear transformation $x\mapsto Ax+b$ with $A={\sf diag}(1.5,1.2,1.6,2,1.8)$ and $b=(1,0,0,1,0)$ to generate covariates on the target domain. As before, we split the data into two parts; we keep $75\%$ of the data as it is, and shift the rest $25\%$ of the data. The results are given in Table 6 and Figure 5. The second column of Table 6, namely Our Method (Without OT) is the aggregation method proposed in this paper, without domain alignment through optimal transport, i.e., we use source data to construct the prediction interval and use the same prediction interval on the target domain. As evident from Table 6, our method outperforms all other methods in terms of the average width of the prediction interval while maintaining a good coverage guarantee.
Airfoil data
Outcome Our Method Our Method (without OT) WVAC WQC
Coverage (Median) 0.928 0.749 0.984 0.952
Coverage (IQR) 0.035 0.22 0.024 0.077
Bandwidth (Median) 15.075 18.512 36.298 32.143
Bandwidth (IQR) 1.638 3.089 10.619 8.364
Bandwidth (Median for Coverage $>$ 95%) 16.429 25.268 37.783 36.433
Table 6: Experimental results for the airfoil data using optimal transport
(a) Our-Width
(b) Our-Coverage
(c) WQC-Width
(d) WQC-Coverage
(e) WVAC-Width
(f) WVAC-Coverage
Figure 3: Experiments on real estate data
(g) Our-Width
(h) Our-Coverage
(i) WQC-Width
(j) WQC-Coverage
(k) WVAC-Width
(l) WVAC-Coverage
Figure 4: Experiments on Energy efficiency data
(a) Our-Width
(b) Our-Coverage
(c) Our(No OT)-Width
(d) Our(No OT)-Coverage
(e) WQC-Width
(f) WQC-Coverage
(g) WVAC-Width
(h) WVAC-Coverage
Figure 5: Experiments on Airfoil data using optimal transport